OpenStudy (anonymous):

WHAT EVER U WANT ILL GIVE NEED HELP Samantha puts an initial $500 into a savings account. The account has a 4% annual compound interest rate. What is the function that best represents Samantha's account and how much does she have in savings after 5 years? Round your answer to the nearest dollar. Hint: Use the formula, f(x) = P(1 + r)x. f(x) = 500(1.4)x, $2,689 f(x) = 500(0.04)x, $512 f(x) = 500(1.04)x, $608 f(x) = 500(1.04)x, $527

OpenStudy (anonymous):

@nonon

OpenStudy (anonymous):

In the formula: r is the rate, P is the initial amount, x is the number of years so just plug in 0.04 for r and 500 for P. f(x) = 500(1.04)^x plug in 5 for x $608 phew

OpenStudy (anonymous):

I want a teddy bear

OpenStudy (anonymous):

THANK U FOLLOW ME AT JHERRERA (INSTAGRAM) AND IM SRY I CANT GIVE U A TEDDY BEAR XD

OpenStudy (anonymous):

:,(

OpenStudy (anonymous):

can U help me with another one @nonon

OpenStudy (anonymous):

suuure

OpenStudy (anonymous):

The sides of a square are three to the power of two sevenths inches long. What is the area of the square? nine to the power of four sevenths square inches three to the power of four sevenths square inches three to the power of the fraction four over forty nine square inches nine to the power of the fraction four over forty nine square inches

OpenStudy (anonymous):

3^(2/7) ? or (3^2)/7?

OpenStudy (anonymous):

here is the screenshot

OpenStudy (anonymous):

@nonon

OpenStudy (anonymous):

SORRY I ALMOST DOSED OFF WAIT A SECOND

OpenStudy (anonymous):

ok ok um if the bases are the same you can add powers think of it as 3^(2/7) * 3^(2/7)

OpenStudy (anonymous):

so yeah 3^(2/7 + 2/7) is the answer

OpenStudy (anonymous):

these are the answers nine to the power of four sevenths square inches three to the power of four sevenths square inches three to the power of the fraction four over forty nine square inches nine to the power of the fraction four over forty nine square inches

OpenStudy (anonymous):

yeah add them

OpenStudy (anonymous):

B

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

:D teddy bear

OpenStudy (anonymous):

btw realize you can just plug it in the calculator or even google and see which ones match

OpenStudy (anonymous):

sry im acting retartded

OpenStudy (anonymous):

Naaaah just helping you in case you have trouble outside of internet :P

OpenStudy (anonymous):

you got instagram

OpenStudy (anonymous):

I don't

OpenStudy (anonymous):

OMG SO SRY ONE MORE PLEASE Given the function f(x) = 2(3)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points)

OpenStudy (anonymous):

wait is this physics???

OpenStudy (anonymous):

Idk im in 8th but i take 9 and 10th grade math

OpenStudy (anonymous):

wait are there any powers in the equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I'm so confused do you have screenshot?

OpenStudy (anonymous):

oh x is a power of 3

OpenStudy (anonymous):

you meant 2(3)^x

OpenStudy (anonymous):

partA: sectionA: 6-2=4 sectionB: 54-18=36 part B: 36/4 = 9 9 times greater

OpenStudy (anonymous):

yeah you just had to plug in each number for x than find the difference

OpenStudy (anonymous):

Thanks jaredhm29 that was great algebra review (not being sarcastic) :D

OpenStudy (anonymous):

THANK U SO MUCH U WERE SOOOOOOO HELPFUL